PREX Meeting Notes, Jan 11, 2010
Present: Bob Michaels, Ole Hansen, Alexandre Camsonne, John LeRose,
Zafar Ahmed, Rupesh Silwal, Luis Mercado, Paul Souder,
Kent Paschke, Abdurahim Rakham, Richard Wilson, Piotr Decowski,
Mark Dalton, Gregg Franklin, Brian Quinn, Megan Friend,
Paul Souder, Guido Urciuoli, Dustin McNulty
Reminder list of shutdown activities.
Be Plug problem (Bob)
- vendor made a mistake, gave CuBe instead of Be.
- It is 97.8% copper. Won't work.
- Investigated substituting carbon (graphite).
- Looks like it may work, trying to get it done in time.
Compton (Abdurahim, Alex)
- New laser hut. General cleanup and replaced HEPA Filter
- Will remove old cavity, start laser hut buildup this week
- Survey of photon line
- Edet gets removed this week
- put in detector group lab
- will think of electronics improvements (other amps, etc)
- replace 1st plane
- improve shielding
- put back on beamline last week Feb
GEM Detectors (Ole)
- Nilanga has 2 pairs: enough for 1 HRS
- Can use old RICH DAQ.
- New advanced electronics from Italy might be ready on time.
- Tracking software exists, needs a new decoder.
- Shift signup available: HERE. Please sign up for shifts !
- Dustin McNulty
- SU scintillators for Moller (with Zafar)
- Systematics of Qsq weighting due to detector pulse heights.
- Gregg Franklin
- improving Compton Monte Carlo
- Megan Friend
- energy dependent asymmmetry analysis
- Was requested to help with comparisons to SU event-trig analysis
- Piotr Decowski
- PREX detector design complete
- all quartz ordered
- testing with 14 cm quartz block
- Zafar Ahmed
- pulser tests of FADC DAQ for Moller
- working with Dustin on SU scint. for Moller
- John LeRose
- awaiting functioning septum magnet; will do field map
- checked field probes with Mindy Jen.
- Paul Souder
- Requested that we do more comparisons of SU Compton analysis
to the CMU FADC integrating analysis from Fall 2009 runs.
I think that was it. See you next week !!
Index of Meetings